The graph of the equation below is a circle. What is the length of the radius of the circle?
(x - 6)2 + (y + 5)2 = 152

Answer by stanbon(57347)

What is the length of the radius of the circle?
(x - 6)^2 + (y + 5)^2 = 152
---------------------------
radius = sqrt(156) = sqrt(4*29) = 2sqrt(29)
===============================================
Cheers,
Stan H.